# Exact Chiral Symmetries of 3+1D Hamiltonian Lattice Fermions

### v2

@@ IV Time Reversal Symmetric Single Weyl Fermion in 3+1d
So far, we have considered time-reversal breaking models. We can also construct time-reversal invariant models, at the cost of making the symmetry generator slightly more not-on-site. As long as $S(\textbf{k})$ is a smooth function of the momentum, then the charge density in real space will be a sum of terms with faster-than-polynomial decay. Such “almost-local” operators share many properties with local operators, while being closed under Hamiltonian evolution generated by such terms [26, 27]. This will allow us to employ bump functions and partitions of unity in momentum space.
@@
V Parity anomaly of a Single Dirac fermion with time-reversal symmetry in 2+1d
@@
VI Discussion
The chiral symmetry in 3+1d has a nearest-neighbor charge operator, but is non-quantized, as it must be by the no-go theorem of [19] and our arguments above. The two $SU(2)$ generators we constructed are quantized (and one is on-site), but do not satisfy the expected $SU(2)$ Lie algebra, instead generating an infinite-dimensional Onsager algebra. This is consistent with the anomaly since either $U(1)$ symmetry on its own is anomaly-free. In this way, it is structurally similar to the $U(1)_{V}\times U(1)_{A}$ anomaly realized in 1+1D in [24, 25], where both generators are quantized and one is on-site, but they don’t commute.
There is a fun example which avoids this no-go argument in 2+1d. Let us take a 2d square lattice of spin-$\frac{1}{2}$ degrees of freedom, and take $\hat{Q}$ to be the toric code Hamiltonian [28], which has an integer spectrum. It also commutes with time reversal given by complex conjugation, and together, these two symmetries generate a bosonic $U(1)\rtimes T$ parity anomaly [29]. The ground state of the toric code Hamiltonian is indeed non-trivial, and sufficient to match this parity anomaly.
Finally, we would eventually like to make lattice models of chiral gauge theories. Although we have produced models with chiral symmetries, the not-on-siteness of the symmetries makes them difficult to gauge. In fact one expects on general grounds that because of the nonvanishing ’t Hooft anomalies, we should not be able to gauge these symmetries. For interesting anomaly cancellation, such as in the Standard Model, one needs low energy fermions of different charges. It is not clear models of the kind we studied can produce such charge assigments. It would be very interesting if in some more complicated models with not-on-site chiral symmetries having vanishing ’t Hooft anomalies, if we can nonetheless find some way to gauge them, as we done for discrete symmetries in 1+1D in [30].

### v3

@@ Appendix C Time Reversal Symmetric Single Weyl Fermion in 3+1D
As long as $S(\textbf{k})$ is a smooth function of the momentum, then the charge density in real space will be a sum of terms with faster-than-polynomial decay. Such ``almost-local'' operators share many properties with local operators. In particular, time evolution under almost-local operators is locality-preserving, in the sense of having a Lieb-Robinson bound [34], mapping almost-local operators to almost-local operators.
Allowing these almost-local operators will allow us to employ bump functions and partitions of unity in momentum space and greatly simplifies the constructions.
@@
Appendix D Parity anomaly of a Single Dirac fermion with time-reversal symmetry in 2+1D
@@
Appendix E Locality of the Spectral Projector
Suppose $h(\textbf{k})$ describes a single-particle Hamiltonian which is exponentially decaying in real space, with a zero-energy band and all other bands occur outside the energy window (1,1)(-1,1). We consider the ``Riesz projector''
$\begin{gathered}P(\textbf{k})=\frac{1}{2\pi i}\oint_{C}\frac{dz}{z-h(\textbf{k})},\end{gathered}$ (28)
where $C$ is a positively-oriented circle of radius <1<1 centered around the origin. Going to the eigenbasis of $h(\textbf{k})$ and evaluating by residues, this gives the projector we discussed in the main text (with $h(\textbf{k})=S_{\text{chiral}}(\textbf{k})$).
Consider analytically continuing this expression to complex k. Since $h(\textbf{k})$ corresponds to an exponentially decaying Hamiltonian in real space, it is analytic in some strip $|{\rm Im}(k_{i})|<\epsilon$. Also in this strip, the zeros of $z-h(\textbf{k})$ move continuously in the $z$ plane, so we can choose some sub-strip $|{\rm Im}(k_{i})|<\epsilon_{1}\leq\epsilon$ where no zero crosses our small circle $C$. Therefore, in that strip $P(\textbf{k})$ will be analytic and so $P(\textbf{k})$ is exponentially-decaying in real space.